Determine if HNO3 can dissolve in the following metal:

1) Write a balanced chemical reaction showing how the 2.09g Cu dissolves in HNO3.

I had put down: Cu(s) + 4HNO3(aq) ——> Cu(NO3)2(aq) + 2NO2(g) + 2H2O(l)
but it's wrong.

2) Determine the minimum volume of 6.3M HNO3 required to completely dissolve the sample (in mL).

Thank you!!
1
Expert's answer
2013-04-10T10:45:07-0400
The reaction of copper metal with a dilute nitric acid will take place according to the following equation:
3Cu(s) + 8HNO3(aq) ——> 3Cu(NO3)(aq) + 2NO(g) + 4H2O(l)

To calculate the amount of material we need copper:

n(Cu) = 2.09 g / 64 g/mol
n(Cu) = 0.0326 mol
n(NO) = (n(Cu)*3) / 2
n(NO) = 0.049 mol

The quantity of substance with nitric acid, we can calculate on the basis of the following:

2) n(HNO3) = 8n(Cu)/3
n(HNO3) = 8*0.0326 mol/3 = 0.087 mol
V = Cm*n
V(HNO3) = 6.3 mol/L * 0.087 mol = 0.547 L
V(HNO3) = 547 mL
Answer:
V(HNO3) = 547 mL